OpenStudy (thatawayz):

An observer (O) is located 900 feet from a building (B). The observer notices a helicopter (H) flying at a 49° angle of elevation from his line of sight. How high is the helicopter flying over the building?

OpenStudy (thatawayz):



OpenStudy (kkutie7):



OpenStudy (kkutie7):

SOHCAHTOA you can use one of the following to find your answer: \[sin(49)=\frac{o}{h}=\frac{h}{hypotenuse}\] \[cos(49)=\frac{a}{h}=\frac{900ft}{hypotenuse}\] \[tan(49)=\frac{o}{a}=\frac{h}{900ft}\] use one of these to calculate the height. Good luck